Make New Entry,
Make Followup Entry
User name Sasha
Log entry time 20:09:25 on March 03, 2012
Entry number 363148
Preliminary Moller results on March 03 14:00 - 19:40
- Crew: Roman, Sasha, Vadym
- Beam energy=2253.65MeV
- The raster was OFF.
- Beam tuning took ~1 hour.
- Helicity frequency=960Hz, delay=8windows, pattern=quartet.
- FADC DAQ was not running (most likely CPU is dead again).
- 14:55-15:34 (Runs 15193-15201) detector HV tuning.
- Runs 15202-15207 beam polarization measurement. Target 2 was used.
- Preliminary beam polarization is -79.73%
- Runs 15207-15210 - BCM calibration (Target 2)
- Runs 15211-15220 - Q1 quad scan (good agreement with GEANT
- Runs 15221-15229 - Q3 quad scan (good agreement with GEANT
- Runs 15230-15231 - Beam polarization measurement
- Runs 15232-15237 - Target 4 at 4A BOP(low magnetic field)
- Runs 15238-15241 - BCM calibration (Target 3)
- Runs 15242-15247 beam charge asymmetry test.
- Set 1: -0.0092. Set2: -0.53*10-2. Set3: -0.9*10-3.
- 20:00: Moller target is parked, Moller target movement is
unmasked. Moller dipole is OFF, quads settings are restored.
A copy of this log entry has been emailed to: firstname.lastname@example.org, email@example.com, firstname.lastname@example.org